# DFT+DMFT study of correlated electronic structure in the monolayer-trilayer phase of La3Ni2O7

### v2

@@ Abstract
By preforming DFT+DMFT calculations, we systematically investigate the correlated electronic structure in the newly discovered monolayer-trilayer (ML-TL) phase of La3Ni2O7 (1313-La327). Our calculated Fermi surfaces are in good agreement with the result of angle-resolved photoemission spectroscopy. We find that 1313-La327 is a multiorbital correlated metal. An unstable orbital-selective Mott behavior is found in ML. The ML Ni-3 orbital exhibits a Mott behavior, while the ML Ni-3 orbital is metallic due to self-doping. Quasi-particle peaks appear at the Fermi level due to an interorbital hopping between the localized ML Ni-3 and delocalized Ni-3 orbitals. We also find a large static local spin susceptibility of ML Ni, suggesting that there is large spin fluctuation in 1313-La327. The TL Ni- orbitals possess similar electronic correlation to those in La4Ni3O10. The orbitals of the outer-layer Ni in TL show non-Fermi liquid behaviors. Besides, large weight of high-spin states are found in TL-outer Ni and ML Ni, implying Hundness. Under 16 GPa, a Lifshitz transition is revealed by our calculations and a La-related band crosses the Fermi level. Our work provides a theoretical reference for studying other potential mixed-stacked nickelate superconductors.
@@ I Introduction
In this paper, we perform the density functional theory plus dynamical mean-field theory (DFT+DFMT) calculations to study the correlated electronic structure in 1313-La327. Our calculated Fermi surfaces are in good agreement with the angle-resolved photoemission spectroscopy (ARPES) results. We find that 1313-La327 is a multiorbital correlated metal and a Lifshitz transition is observed under 16 GPa. Furthermore, our results suggest that 1313-La327 is a unique platform, which may include Mott and Hund physics. Specifically, the calculated self-energy functions and orbital-resolved spectral functions show that the ML Ni-3 orbitals exhibit a Mott behavior, while the ML Ni-3 orbitals are metallic due to the self-doping between the ML and TL. An unstable orbital-selective Mott phase (OSMP) is found in the ML. Besides, we find large weights of high-spin states in TL-outer Ni and ML Ni, which indicates Hundness. The calculated static local spin susceptibility suggests that there is a complicated magnetic property in 1313-La327.
@@ II.1 Electronic structure
In order to make clear the characters of bands and the renormalization effect, we show the DFT calculated band structure and orbital-projected density of states (DOS) of 1313-La327 under ambient and 16 GPa pressures in Figs. 1(b-d) and (f-h). The DFT results suggest that 1313-La327 is a multiorbital metal, where the orbitals of all three kinds of Ni atoms contribute the bands near the Fermi level. And the bandwidth of the Ni- orbitals is 4 eV, which is consistent with other previous theoretical reports [27, 31]. Furthermore, we find that at ambient pressure La-5 orbitals contribute little to the DOS around the Fermi level [Fig. 1(c)]. However, under 16 GPa pressure the DOS of the La-5 orbitals becomes finite at the Fermi level [Fig. 1(g)], which is mainly contributed by the surface La atoms of the ML and TL.
In Figs. 2(b) and (c), we show the DFT+DMFT calculated Fermi surfaces at = 0 plane under ambient pressure and 16 GPa pressure, respectively. And the ARPES Fermi surfaces under ambient pressure [30] are exhibited in Fig. 2(a). Our calculated Fermi surfaces under ambient pressure are in good agreement with the ARPES results. Besides the , , and pockets which are similar to those in 2222-La327, a newly discovered pocket in 1313-La327 has also been well reproduced in our theoretical calculations. This pocket shows a diamond-like pattern around the point in Fig. 2(b). However, in Fig. 2(c), under 16 GPa pressure, it is clearly seen that there is a ring-like Fermi surface with large quasi-particle weight, which is quite close to the pocket. This is the La-related band, which crosses the Fermi level around the point. This phenomenon has not been found in our previous studies about 2222-La327 [14] or La4310 [24]. For example, the La-related bands in 2222-La327 have no contribution to the Fermi surfaces. But it is reminiscent of the self-doping in infinite-layer nickelates La1-xSrxNiO2, where La-related band crosses the Fermi level [32, 33].
@@ II.5 Results for low and zero temperatures
In Figs. 4(c) and (f), we show the spectral functions of the ML Ni- orbitals at zero temperature, where the quasi-particle resonance peaks at the Fermi level remain. And this situation is well consistent to an unstable OSMP scenario. Specifically, the ML Ni-3 and Ni-3 orbitals would be insulating and metallic, respectively, if there was no interorbital hopping. The appearance of quasi-particle peaks at the Fermi level is attributed to the finite interorbital hopping between the ML Ni-3 and Ni-3 orbitals. And a previous theoretical report regarding 2222-La327 suggests that there is a finite effective interorbital hopping between the Ni-3 and Ni-3 orbitals [38]. The unstable OSMP in 1313-La327 is a realistic example of theoretical report [39]. A Mott insulating (heavy) orbital is unstable when there is a finite interorbital hopping to a different metallic (light) orbital. And the of the heavy orbital develops a thin quasi-particle peak at the Fermi level. This unstable OSMP scenario is reminiscent of Kondo physics where hybridization happens between local moments and conducting electrons. But here the hybridization is between local moments (here ML Ni-3 orbitals) and a non-Fermi liquid (here ML Ni-3 orbitals).
@@ III Discussion and Conclusion
(ii) The interorbital hopping between the localized ML Ni-3 and delocalized Ni-3 orbitals breaks the OSMP of the ML, giving rise to quasi-particle peaks in the spectra of the ML Ni- orbitals at the Fermi level. This suggests that 1313-La327 is a realistic example of the unstable OSMP scenario [39].
@@
In Summary, we have systematically studied the correlated electronic structure in the newly discovered 1313-La327. Our calculated Fermi surfaces are in good agreement with the ARPES results. We find that 1313-La327 is a multiorbital correlated metal. Due to the self-doping from the ML to TL, the ML Ni-3 orbitals are metallic and show a non-Fermi liquid behavior. But the ML Ni-3 orbitals show a Mott behavior. And the interorbital hopping between the localized ML Ni-3 and delocalized Ni-3 orbitals causes an unstable OSMP. Our calculations about the static local spin susceptibility suggest that there may be large spin fluctuation in ML Ni and 1313-La327 may have a more complicated magnetic property. As for the TL in 1313-La327, the TL-outer Ni- orbitals show a non-Fermi liquid behavior, which is similar to the case of La4310. The large weights of high-spin states of ML Ni and TL-outer Ni indicate Hundness in 1313-La327. Under 16 GPa pressure, a Lifshitz transition is observed and a La-related band crosses the Fermi level. Our studies give a new perspective to understand this ML-TL stacking structure and provide a theoretical reference for studying other potential mixed-stacked nickelate superconductors.

### v3

@@ Abstract
By preforming DFT+DMFT calculations, we systematically investigate the correlated electronic structure in the newly discovered monolayer-trilayer (ML-TL) phase of La3Ni2O7 (1313-La327). Our calculated Fermi surfaces are in good agreement with the result of angle-resolved photoemission spectroscopy. We find that 1313-La327 is a multiorbital correlated metal. An orbital-selective Mott behavior is found in ML in our zero- and finite-temperature calculations. The ML Ni-3 orbital exhibits a Mott behavior, while the ML Ni-3 orbital is metallic due to self-doping. We also find a large static local spin susceptibility of ML Ni, suggesting that there is large spin fluctuation in 1313-La327. The TL Ni- orbitals possess similar electronic correlation to those in La4Ni3O10. The orbitals of the outer-layer Ni in TL show non-Fermi liquid behaviors. Besides, large weight of high-spin states are found in TL-outer Ni and ML Ni, implying Hundness. Under 16 GPa, a Lifshitz transition is revealed by our calculations and a La-related band crosses the Fermi level. Our work provides a theoretical reference for studying other potential mixed-stacked nickelate superconductors.
@@ I Introduction
In this paper, we perform the density functional theory plus dynamical mean-field theory (DFT+DFMT) calculations to study the correlated electronic structure in 1313-La327. Our calculated Fermi surfaces are in good agreement with the angle-resolved photoemission spectroscopy (ARPES) results. We find that 1313-La327 is a multiorbital correlated metal and a Lifshitz transition is observed under 16 GPa. Furthermore, our results suggest that 1313-La327 is a unique platform, which may include Mott and Hund physics. Specifically, the calculated self-energy functions and orbital-resolved spectral functions show that the ML Ni-3 orbitals exhibit a Mott behavior, while the ML Ni-3 orbitals are metallic due to the self-doping between the ML and TL. An orbital-selective Mott phase (OSMP) is found in the ML. Besides, we find large weights of high-spin states in TL-outer Ni and ML Ni, which indicates Hundness. The calculated static local spin susceptibility suggests that there is a complicated magnetic property in 1313-La327.
@@ II.1 Electronic structure
In order to make clear the characters of bands and the renormalization effect, we show the DFT calculated band structure and orbital-projected density of states (DOS) of 1313-La327 under ambient and 16 GPa pressures in Figs. 1(b-d) and (f-h). The DFT results suggest that 1313-La327 is a multiorbital metal, where the orbitals of all three kinds of Ni atoms contribute the bands near the Fermi level. And the bandwidth of the Ni- orbitals is 4 eV, which is consistent with other previous theoretical reports [27, 31]. Furthermore, we find that at ambient pressure La-5 orbitals contribute little to the DOS around the Fermi level [Fig. 1(c)]. However, under 16 GPa pressure the DOS of the La-5 orbitals becomes finite at the Fermi level [Fig. 1(g)]. And a La-related band crosses the Fermi level around the point, which is mainly contributed by the surface La atoms of the ML and TL.
In Figs. 2(b) and (c), we show the DFT+DMFT calculated Fermi surfaces at = 0 plane under ambient pressure and 16 GPa pressure, respectively. And the ARPES Fermi surfaces under ambient pressure [30] are exhibited in Fig. 2(a). Our calculated Fermi surfaces under ambient pressure are in good agreement with the ARPES results. Besides the , , and pockets which are similar to those in 2222-La327, a newly discovered pocket in 1313-La327 has also been well reproduced in our theoretical calculations. This pocket shows a diamond-like pattern around the point in Fig. 2(b). However, in Fig. 2(c), under 16 GPa pressure, it is clearly seen that there is a ring-like Fermi surface with large quasi-particle weight, which is quite close to the pocket. This phenomenon has not been found in our previous studies about 2222-La327 [14] or La4310 [24]. For example, the La-related bands in 2222-La327 have no contribution to the Fermi surfaces. But it is reminiscent of the self-doping in infinite-layer nickelates La1-xSrxNiO2, where La-related band crosses the Fermi level [32, 33].
@@ II.5 Results for low and zero temperatures
In Figs. 4(c) and (f), we show the spectral functions of the ML Ni- orbitals at zero temperature, where the quasi-particle resonance peaks at the Fermi level remain. And this situation is well consistent to an OSMP scenario. Specifically, the ML Ni-3 and Ni-3 orbitals would be insulating and metallic, respectively, if there was no interorbital hopping. The appearance of quasi-particle peaks at the Fermi level is attributed to the finite interorbital hopping between the ML Ni-3 and Ni-3 orbitals. And a previous theoretical report regarding 2222-La327 suggests that there is a finite effective interorbital hopping between the Ni-3 and Ni-3 orbitals [38]. The OSMP in 1313-La327 is a realistic example of theoretical report [39]. The of the heavy orbital develops a thin quasi-particle peak at the Fermi level. This OSMP scenario is reminiscent of Kondo physics where hybridization happens between local moments and conducting electrons. But here the hybridization is between local moments (here ML Ni-3 orbitals) and a non-Fermi liquid (here ML Ni-3 orbitals).
@@ III Discussion and Conclusion
(ii) The interorbital hopping between the localized ML Ni-3 and delocalized Ni-3 orbitals breaks the OSMP of the ML, giving rise to quasi-particle peaks in the spectra of the ML Ni- orbitals at the Fermi level. This suggests that 1313-La327 is a realistic example of the OSMP scenario [39].
@@
In Summary, we have systematically studied the correlated electronic structure in the newly discovered 1313-La327. Our calculated Fermi surfaces are in good agreement with the ARPES results. We find that 1313-La327 is a multiorbital correlated metal. Due to the self-doping from the ML to TL, the ML Ni-3 orbitals are metallic and show a non-Fermi liquid behavior. But the ML Ni-3 orbitals show a Mott behavior. And the interorbital hopping between the localized ML Ni-3 and delocalized Ni-3 orbitals causes an OSMP. Our calculations about the static local spin susceptibility suggest that there may be large spin fluctuation in ML Ni and 1313-La327 may have a more complicated magnetic property. As for the TL in 1313-La327, the TL-outer Ni- orbitals show a non-Fermi liquid behavior, which is similar to the case of La4310. The large weights of high-spin states of ML Ni and TL-outer Ni indicate Hundness in 1313-La327. Under 16 GPa pressure, a Lifshitz transition is observed and a La-related band crosses the Fermi level. Our studies give a new perspective to understand this ML-TL stacking structure and provide a theoretical reference for studying other potential mixed-stacked nickelate superconductors.